OpenStudy (anonymous):

How do I solve the equation 2t/3=4/9

OpenStudy (anonymous):

If the equation is:\[\frac{ 2t }{ 3}=\frac{ 4 }{ 9 }\] and you're solving for t, just move everything from the left to the right side by multiplying the 3 from the left, and then dividing the 2 from the right-> \[2t=\frac{ 3\times4}{9 }\]\[t =\frac{ 3\times4 }{ 9\times2 }=\frac{ 12 }{ 18 }=\frac{ 2 }{ 3}\] so t= 2/3